OpenStudy (anonymous):

Can someone teach me how to factor a perfect square? I have an example in my notes and I know what to do until I get to the perfect square. I appreciate it.

OpenStudy (anonymous):

sure, do you have an example you could provide?

OpenStudy (anonymous):

Convert the following functions to standard form. These are all quadratic functions Y=-4(x+2)(x+7)

OpenStudy (anonymous):

I went through (hopefully correctly) and I'm stuck at y=-4(x^2-36x+1)-56

OpenStudy (anonymous):

The example from my notes confuses me a bit

OpenStudy (triciaal):

to factor a perfect square you are just using the positive and the negative of the same number example 16 sq rt of 16 4 but it could also be -4 because when you multiply a negative by a negative the result is positive +4*+4 = +16 and -4*-4 = + 16 144 = 12^2 but it is also -12^2

OpenStudy (triciaal):

any questions before i look at your notes?

OpenStudy (triciaal):

ok so I looked at your notes and what I got from it is you don't need to factor a perfect square you need to rewrite the quadratic so you can square the factor.